### v1

@@ 2.1 Non-cooperate Game
$\hat{v}_{t}^{i}$ is a measurable function w.r.t $(t,x)\in[0,T]\times\mathbb{R}$, and satisfies the terminal condition $K_{i}(T,\cdot)=\Psi(\cdot)$. From the proof of subsection 2.1, $a(t)$, $B(t)$ and $F_{i}(t)$ are bounded on $[0,T]$, so $K_{i}(t,x)$ satisfies the quadric growth condition (i.e. there exists $C>0$ such that $|K_{i}(t,x)|\leq C(1+x^{2}),\ \forall(t,x)\in[0,T]\times\mathbb{R}$). Besides, (2.10)-(2.12) imply that, (2.9) solves the equation
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}$
$=\frac{\partial{K}_{i}(t,x)}{\partial t}+\hat{v}_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({\hat{v}_{t}^{i}})^{2}+\hat{v}_{t}^{i}\bar{P}(t)+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
The result then follows from Pham (2009). ∎
Now, we will approximate the average commodity process with a deterministic process. The underlying reason for this approximation can be explained as follows: by denoting the average of $X_{t}^{i}$ by $\overline{X}_{t}$, the average of this optimal control has the form $\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c}$, which leads to the mean commodity process evolving as follows
@@
It is worth noticing that $\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i}$ (the average of the independent Brownian motions) forms a Brownian motion with a small magnitude, denoted by $\sigma dB_{t}$, where $\sigma:=\frac{1}{N}\sqrt{\sum_{i=1}^{N}\sigma_{i}^{2}}$ represents its magnitude and $B_{t}$ can be interpreted as the standardized form of the Brownian motion obtained by aggregating the independent noise. It is easily checked that this magnitude gradually decreases to zero as $N$ increases, provided that $\{\sigma_{i}\}_{i=1}^{N}$ is bounded. Therefore, by using a mean-field approximation, among which we let $N\to\infty$, we will approximate the average commodity process with a determinist function $\bar{x}(t)$, whose dynamics read
@@
By consistency condition, the approximate average trading rate process it induces should satisfy
$\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}=\bar{Q}_{t},\quad t\in[0,T].$
Using the relationship between $\bar{P}(t),B(t)$ and $\bar{Q}(t)$, it is equivalent that the following equation system holds
@@
By introducing matrix notation, we obtain an alternative formulation, which requires the following process
@@
to hold true simultaneously with the following initial or terminal conditions
@@
The existence of solutions to (2.13) and (2.14) is crucial, for it determines whether or not the requirement in step 3 can be satisfied. The main difficulty in determining whether this equation has a solution lies in the fact that its boundary values are specified at different points in time. Therefore, when investigating the existence of its solution, we first analyze the properties of solutions given initial values at the same point, which will be presented in subsection 2.1 and subsection 2.1. They can be regarded as a preparation for the subsection 2.1, which proves that, under an assumption to be proposed (i.e. Assumption $\boldsymbol{(A_{1})}$), the existence and uniqueness of solutions to (2.13) and (2.14) hold true.
@@
subsection 2.1 shows that the solution for (2.13) is uniquely determined by $B(0)=b_{0}$ once the conditions $\bar{x}(0)=\bar{x}_{0}^{N}$ and $\bar{P}(0)=p_{0}$ are given. Therefore, it remains to adjust the value of $b_{0}$ so that $B(T)=-\gamma\zeta$ holds true.
To this end, we construct the following process.
@@
Now, we propose Assumption $\boldsymbol{(A_{1})}$, and we claim that under this assumption, (2.13), combined with (2.14), has a solution.
@@
If the unique solution to (2.15) satisfies $B_{1}(T)\neq 0$, (i.e. Assumption $\boldsymbol{(A_{1})}$ holds true), then (2.13), combined with (2.14), gives a unique solution.
We introduce another assumption, which places a constraint on the magnitude of the elements within the sets $\{\sigma_{i}\}_{1\leq i\leq N}$ and $\{x_{0}^{i}\}_{1\leq i\leq N}$.
@@
Assume that Assumption $\boldsymbol{(A_{1})}$ and $\boldsymbol{(A_{2})}$ hold true, and the unique solution to (2.13) and (2.14) is given by (B(t),x¯(t),P¯(t))(B(t),\ $\bar${x}(t),\ $\bar${P}($)^{\top}, t$hen the following system
@@
is an $\epsilon$-Nash equilibrium. Moreover, specific calculations show that $\epsilon_{N}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)}$ in (2.4).
Remark 2.
In its proof, the $O(1)$ numbers between adjacent lines may not be the same. This convention is also applied to the proofs in the cooperate games case. In addition, to present our derivation more clearly, we may not abbreviate specific important $O(1)$ numbers as $O(1)$, but instead denote them with specific letters.
In order to prove Theorem 2.6, we will introduce three lemmas. First, we introduce Lemma 2.1.
@@
If $\hat{v}^{i}$, $\hat{X}^{i}$ and $\hat{Q}$ are given by (2.16), then there exists a positive constant $K^{\prime}$ such that
@@
The proofs of all lemmas in this paper have been left to Appendix A.
To show that the system indeed induces an $\epsilon$-Nash equilibrium, we will show that for any agent $i$, its expected cost $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ does not decrease much if it changes its strategy unilaterally. Specifically, for agent $i(i=1,\ldots,N)$, we consider the case in which it adopts any admissible strategy $v^{i}\in\mathcal{A}$, while the strategies of all other agents remain fixed as $\hat{\boldsymbol{v}}^{-i}$. Then, agent $i$’s commodity evolves as follows
@@
and the evolution process for the price changes to
@@
To prove Theorem 2.6, we will compare the expected cost $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ with $J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$. Our aim is to prove (2.4) with $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$, or equivalently, for $i=1,\ldots,N$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)},\quad\forall\ v^{i}\in\mathcal{A}.$ (2.17)
In order to prove (2.17) , we need coercive condition, whose detailed statement is provided in Lemma 2.3. Prior to that, we present Lemma 2.2, which will facilitate a clear proof of Lemma 2.3.
@@
For $N\in\mathbb{N}^{+}$, we construct the process
@@
We claim that for any given $\epsilon_{1}^{*}>0$, $\epsilon_{2}^{*}>0$ and $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, we have
$\mathbb{E}\left[\int_{0}^{T}v_{t}(\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t})dt\right]\geq 0,\quad\forall v\in\mathcal{A}.$ (2.18)
Next, we introduce the coercive condition.
@@
There exists a constant $C_{0}$, which is independent of $N$, such that the coercive condition
@@
is satisfied for $N>\frac{4(\sqrt{2}+1)}{c}$.
Now, we return to our proof of Theorem 2.6.
@@
For $\forall N>\frac{4(\sqrt{2}+1)}{c}$, if $\frac{c}{4}\|v^{i}\|^{2}+C_{0}\geq K^{\prime}$, then it follows from Lemma 2.3 that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}).$
This will make (2.17) hold. Otherwise, $\|v^{i}\|^{2}$ is bounded by $\frac{4(K^{\prime}-C_{0})}{c}$. Therefore, it suffices to prove (2.17) under the case where $\|v^{i}\|$ is bounded by an $O(1)$ constant.
As mentioned above, $\bar{P}(t)$ is constructed for approximating $P_{t}$ and $\hat{P}_{t}$. This inspires us to write the difference of the expected cost as
@@
To finish the proof, it suffices to show that in the case where $\|v^{i}\|$ is bounded by an $O(1)$ number,
$I_{1}\geq 0,\quad I_{2}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)},\quad I_{3}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)}.$
The first part is easy. We have proved that $\hat{v}_{t}^{i}$ is optimal for minimizing
$\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right].$
This gives $I_{1}\geq 0$.
The second and the third part is based on the fact that, when $\|v^{i}\|$ is bounded by an $O(1)$ number, $|I_{2}|$ and $|I_{3}|$ are bounded by $O(\frac{1}{\sqrt{N}})$ numbers. Specifically, we calculate both the difference between $\hat{P}_{t}$ and $\bar{P}(t)$ and that between $\bar{x}(t)$ and $\overline{\hat{X}}_{t}$. Direct calculation yields
@@
It follows from Gronwall Lemma and Assumption $\boldsymbol{(A_{2})}$ that
@@
which implies (from the Cauchy-Schwartz inequality)
@@
Then, it is convenient and useful for us to compare $P_{t}$ with $\hat{P}_{t}$,
@@
We deduce from the Cauchy-Schwartz inequality ,
$\|P-\hat{P}\|^{2}\leq\frac{1}{N^{2}}\left(\int_{0}^{T}dt\int_{0}^{t}(e^{-\alpha(t-s)}\alpha)^{2}ds\right)\mathbb{E}\left[\int_{0}^{T}(v_{s}^{i}-\hat{v}_{s}^{i})^{2}ds\right]=O\bigg{(}\frac{1}{N^{2}}\bigg{)}.$
The last equality is based on the fact that $\|\hat{v}^{i}\|$ and $\|v^{i}\|$ are all bounded by $O(1)$ numbers. The above estimates imply that
$\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),\quad\|P-\hat{P}\|=O\bigg{(}\frac{1}{N}\bigg{)},\quad\|P-\bar{P}\|\leq\|P-\hat{P}\|+\|\hat{P}-\bar{P}\|\leq O\left(\frac{1}{\sqrt{N}}\right),$
where the last result follows from the triangle inequality. By using the Cauchy-Schwartz inequality again
@@
$|I_{3}|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}-\bar{P}\|=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
We complete the proof for the second and the third part. Finally, we conclude that
@@
2.2 The Cooperate Games
In this subsection, we are looking for a set of strategies that achieve an $\epsilon$-optimal expected social cost. Since our objective is to minimize $J_{soc}$, it is necessary for us to comprehensively measure the impact of agent $i$’s strategy on $J_{soc}$. To this end, We decompose the price process $P$ into components that are influenced and not influenced by agent $i$, respectively, and proceed to analyze the effect of $v^{i}$ on $J_{soc}$.
@@ 2.2 The Cooperate Games
    Show that under the optimal strategies of $J_{soc}$, agent $i$’s strategy must be the minimizer of the following auxiliary control problem:
@@
    We approximate $\frac{1}{N}\sum_{k\neq i}\check{v}_{t}^{k}$ with a continuous function $(\bar{q}(t))_{0\leq t\leq T}$. Then, we derive a continuous function $(\bar{p}(t))_{0\leq t\leq T}$, which can be viewed as the approximation of price process $(P_{t})_{t\in[0,T]}$. It evolves according to
@@
    For $i=1,\ldots,N$, by applying the approximation introduced above, we obtain the following auxiliary objective function for agent $i$ from $\check{J}_{i}$,
@@
    subject to $(X^{i},u^{i})$ with (2.20).
@@
    Solve control problem $J^{\prime}_{i}$, obtain the optimal control strategies, and construct the consistence condition which is satisfied by $\bar{q}$. Finally, we show that the strategies derived from the solution to this consistence condition indeed has asymptotic social optimality.
Our first step is to evaluate precisely how the strategy of a certain agent affects the expected social cost $J_{soc}$. To do this, we divide $J_{soc}$ into three parts, one of them being closely related to its own strategy, the rest being independent of its control. To be more specific, we state that:
@@
If the strategies $\{\check{v}_{t}^{i}\}_{i=1}^{N}$ optimize $J_{soc}$, then for any $i=1,\ldots,N$, $\check{v}_{t}^{i}$ is the solution for minimizing control problem (2.19)-(2.20) when other agents’ strategies are given by $\check{v}^{-i}$.
@@
We divide $J_{soc}$ into different parts. The objective is to evaluate the overall impact of agent $i$’s strategy on $J_{soc}$. Let $J_{soc}=Y_{1}^{i}+Y_{2}^{i}$, for $i=1,\ldots,N$, where
@@
It should be noted that the evolution of $P$ is only partially influenced by strategy $v^{i}$. Therefore, we solve from (2.2)
@@
and apply $Q_{t}=\frac{1}{N}v_{t}^{i}+\frac{1}{N}\sum_{k\neq i}v_{t}^{k}$ to obtain $Y_{2}^{i}=Y_{3}^{i}+Y_{4}^{i}$, where
@@
Observe that $J_{soc}=Y_{1}^{i}+Y_{3}^{i}+Y_{4}^{i}$, and $Y_{1}^{i}$, $Y_{3}^{i}$ are both independent of $v^{i}$. Combining with $Y_{4}^{i}=\frac{1}{N}\check{J}_{i}(v^{i})$, we conclude that if $(\check{v}_{t}^{1},...,\check{v}_{t}^{N})$ minimizes $J_{soc}$, then $\check{v}_{t}^{i}$ minimizes $\check{J}_{i}(v^{i})$ for $i=1,\ldots,N$. The proof is complete. ∎
Subsequently, we solve the optimal control problem (2.22). It follows from (2.20) that
@@
$V_{i}(t,x,u)=\inf_{v^{i}\in\mathcal{A}}\mathbb{E}\left[\int_{t}^{T}\Big{(}\frac{\eta}{2}(X_{s}^{i}-\kappa)^{2}+\frac{c}{2}(v_{s}^{i})^{2}-\bar{q}(s)u_{s}^{i}+v_{s}^{i}\bar{p}(s)\Big{)}ds+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
@@
$\frac{\partial V_{i}}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{\frac{\partial V_{i}}{\partial x}v_{t}^{i}+\frac{\partial V_{i}}{\partial u}(\alpha v_{t}^{i}-\alpha u)+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)\right\}-\bar{q}(t)u+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}V_{i}}{\partial^{2}x}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
and the terminal condition $V_{i}(T,x,u)=\frac{\gamma}{2}(x-\zeta)^{2}$. We consider the ansatz
$V_{i}(t,x,u)=a(t)x^{2}+b(t)x+l(t)u+f_{i}(t).$
The HJB equation then reduces to
$a^{\prime}(t)x^{2}+b^{\prime}(t)x+f_{i}^{\prime}(t)+l^{\prime}(t)u+\inf_{v^{i}\in\mathcal{A}}\Big{\{}(2a(t)x+b(t))v_{t}^{i}+\alpha l(t)v_{t}^{i}-\alpha ul(t)$

### v2

@@ 3 Non-cooperate Game
$\hat{v}:(t,x)\to-\frac{\bar{P}(t)+2a(t)x+B(t)}{c}$ is Lipschitz-continuous w.r.t $(t,x)\in[0,T]\times\mathbb{R}$, implying that (2.1) admits a unique strong solution (Theorem 8.3 in Le Gall, J. F. (2016)). Since $a(t)$, $B(t)$ and $F_{i}(t)$ are bounded on $[0,T]$, the function $K_{i}(t,x)$ satisfies the quadric growth condition; that is, there exists $C>0$ such that $|K_{i}(t,x)|\leq C(1+x^{2}),\ \forall(t,x)\in[0,T]\times\mathbb{R}$. Additionally, the terminal condition $K_{i}(T,\cdot)=\Psi(\cdot)$ is fulfilled. Under these conditions on $K_{i}$, together with
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}_{i}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}$
$=\frac{\partial{K}_{i}(t,x)}{\partial t}+\hat{v}_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({\hat{v}_{t}^{i}})^{2}+\hat{v}_{t}^{i}\bar{P}(t)+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$
we obtain the result by applying the arguments in Theorem 3.5.2, Pham (2009). ∎
We now approximate the average commodity process by a deterministic function. The rationale behind this approximation is as follows. Let $\overline{X}_{t}$ denote the average of $X_{t}^{i}$. Then the average optimal control takes the form
$\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c},$
which implies that the mean trading process evolves as follows
@@
Note that $\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i}$ represents a Brownian motion with magnitude $\frac{1}{N}\sqrt{\sum_{i=1}^{N}\sigma_{i}^{2}}$. This magnitude clearly vanishes to zero as $N\to\infty$ provided that $\sigma_{i}\ (\forall i=1,2,\ldots,N)$ are bounded by a constant independent of $N$. Hence, under the standard mean-field approximation, we replace the average stochastic process by a deterministic function $\bar{x}(t)$, whose dynamics satisfy
@@
By consistency condition, the expectation of the induced average trading process must satisfy
$\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}=\bar{Q}_{t},\quad\forall t\in[0,T].$
Using the relationship between $\bar{P}(t),B(t)$ and $\bar{Q}(t)$, this requirement is equivalent to the following system:
@@
By introducing matrix notation, we obtain an alternative form, which requires the following function system
@@
to hold simultaneously with the following initial or terminal conditions
@@
The existence of solutions to (3.9) and (3.10) is essential, as it determines whether the consistency requirement in Step 3 can be fulfilled. The main challenge here is: the boundary conditions are imposed at different time points.
To tackle this issue, we first study solutions with initial conditions specified at the same time point, and establish their properties in section 3 and section 3. These results serve as preparatory steps for section 3, where we prove that, under Assumption $\boldsymbol{(A_{1})}$ stated below, the coupled system (3.9)–(3.10) admits a unique solution.
@@
section 3 shows that once the conditions $\bar{x}(0)=\bar{x}_{0}^{N}$ and $\bar{P}(0)=p_{0}$ are given, the solution to (3.9) is uniquely determined by the choice of $B(0)=b_{0}$ . Therefore, our remaining task is to adjust the value of $b_{0}$ so that $B(T)=-\gamma\zeta$ holds.
To this end, we construct the following auxiliary function system,
@@
We now introduce Assumption $\boldsymbol{(A_{1})}$ and show that, under this assumption, the coupled system (3.9)–(3.10) indeed admits a solution.
@@
If Assumption $\boldsymbol{(A_{1})}$ holds, then (3.9), combined with (3.10), gives a unique solution.
We further impose an assumption that constrains the magnitudes of the elements in the sets $\{\sigma_{i}\}_{1\leq i\leq N}$ and $\{x_{0}^{i}\}_{1\leq i\leq N}$.
@@
Assume that Assumption $\boldsymbol{(A_{1})}$ and $\boldsymbol{(A_{2})}$ hold, and the unique solution to (3.9) and (3.10) is given by (B(t),x¯(t),P¯(t))(B(t),\ $\bar${x}(t),\ $\bar${P}($)^{\top}, t$hen the following system
@@
is an $\epsilon$-Nash equilibrium. Moreover, specific calculations show that under the proposed strategies $\{\hat{v}_{t}^{i}\}_{i=1}^{N}$, $\epsilon_{N}=O\Big(\frac{1}{\sqrt{N}}\Big)$ in (2.4).
To prove Theorem 3.6, we will need three auxiliary lemmas. As a first step, in order to derive uniform bounds for the norms of the state processes under the optimal control and for the quantities $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\ \forall i=1,2,\ldots,N$, we introduce Lemma 3.1.
@@
There exists a positive constant $K^{\prime}$, independent of $N$ and $i$, such that
@@
The proofs of all lemmas have been left to Appendix A.
[…]
@@
and the price dynamics become
@@
We aim to compare $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ with $J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ and establish (2.4) with $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$, or equivalently,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg(\frac{1}{\sqrt{N}}\bigg),\quad\forall\ v^{i}\in\mathcal{A}_{i}.$ (3.13)
To establish (3.13), we require a coercivity condition, whose precise formulation is given in Lemma 3.3. Before presenting it, we first introduce Lemma 3.2, which will streamline the proof of the coercivity result.
@@
For any $N\in\mathbb{N}^{+}$ and $v\in\mathcal{A}_{i}$, we construct the process
@@
Then, for any given $\epsilon_{1}^{*}>0$, $\epsilon_{2}^{*}>0$ and $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, we have
$\mathbb{E}\left[\int_{0}^{T}v_{t}(\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t})dt\right]\geq 0.$ (3.14)
@@
There exists a constant $C_{0}$, which is independent of $N$ and $i$, such that the coercive condition
@@
is satisfied for any $N>\frac{4(\sqrt{2}+1)}{c}$.
We now return to our proof of Theorem 3.6.
@@
For all $N>\frac{4(\sqrt{2}+1)}{c}$, if $\frac{c}{4}\|v^{i}\|^{2}+C_{0}\geq K^{\prime}$, then by Lemma 3.3, we immediately obtain that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),$
which yields (3.13). Otherwise, $\|v^{i}\|^{2}$ is bounded by $\frac{4(K^{\prime}-C_{0})}{c}$. Hence, it suffices to establish (3.13) in the case where $\|v^{i}\|$ is bounded by an $O(1)$ constant.
As noted earlier, $\bar{P}(t)$ is constructed for approximating $P_{t}$ and $\hat{P}_{t}$. This motivates us to express the difference of the expected cost as
@@
Thus, to finish the proof, it suffices to show
$I_{1}\geq 0,\quad I_{2}=O\Big(\frac{1}{\sqrt{N}}\Big),\quad I_{3}=O\Big(\frac{1}{\sqrt{N}}\Big).$ (3.15)
Since we have already established that $\hat{v}_{t}^{i}$ is optimal for minimizing
$\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right],$
it follows directly that $I_{1}\geq 0$.
We now address the remaining term in (3.15). Direct calculation yields
@@
Applying Grönwall’s lemma and Assumption $\boldsymbol{(A_{2})}$ yield
@@
which further implies, using the Cauchy–Schwarz inequality,
@@
Then, from (2.2),
@@
From the price process (2.2), we deduce by another application of the Cauchy–Schwarz inequality that
$\|P-\hat{P}\|^{2}\leq\frac{1}{N^{2}}\left(\int_{0}^{T}dt\int_{0}^{t}(e^{-\alpha(t-s)}\alpha)^{2}ds\right)\mathbb{E}\left[\int_{0}^{T}(v_{s}^{i}-\hat{v}_{s}^{i})^{2}ds\right]=O\bigg(\frac{1}{N^{2}}\bigg),$
where the last equality follows from the fact that $\|\hat{v}^{i}\|$ and $\|v^{i}\|$ are all bounded by $O(1)$ numbers. These estimates imply
$\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),\quad\|P-\hat{P}\|=O\bigg(\frac{1}{N}\bigg),\quad\|P-\bar{P}\|\leq\|P-\hat{P}\|+\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),$
where the last step applies the triangle inequality. Applying the Cauchy-Schwartz inequality once more gives
@@
$|I_{3}|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}-\bar{P}\|=O\bigg(\frac{1}{\sqrt{N}}\bigg),$
which establishes (3.15). Finally, combining all the above arguments, we conclude that
@@
4 The Cooperate Games
In this section, we aim to identify a set of strategies that achieve an $\epsilon$-optimal expected social cost. Since our objective is to minimize $J_{soc}$, it is essential to quantify the influence of agent $i$’s control on the value of $J_{soc}$. To this end, we decompose the price process $P$ into components that are affected and unaffected by agent $i$, and then analyze how $v^{i}$ contribute to $J_{soc}$.
@@ 4 The Cooperate Games
    We first show that under the optimal strategies of $J_{soc}$, the control of each agent $i$’ must minimize the following auxiliary control problem:
@@
    We approximate the empirical mean $\frac{1}{N}\sum_{k\neq i}\check{v}_{t}^{k}$ with a continuous function $(\bar{q}(t))_{0\leq t\leq T}$. Subsequently, we construct a continuous approximation function $(\bar{p}(t))_{0\leq t\leq T}$ to the price process $(P_{t})_{t\in[0,T]}$, which evolves according to
@@
    For each agent $i=1,\ldots,N$, applying the above approximation yields the following auxiliary objective function,
@@
    subject to the state dynamics $(X^{i},u^{i})$ given by (4.2).
@@
    By solving the control problem $J^{\prime}_{i}$, we obtain the corresponding optimal strategies and derive the consistency condition that the function $\bar{q}$ must satisfy. Finally, we verify that the strategies generated from the solution of this consistency condition achieve asymptotic social optimality.
Our first step is to evaluate precisely the impact of a single agent’s strategy on the expected social cost $J_{soc}$. To do this, we divide $J_{soc}$ into three parts, one of them depends only on agent $i$’s strategy, while the others are independent of it. Specifically, we state that:
@@
If the strategies $\{\check{v}_{t}^{i}\}_{i=1}^{N}$ optimize $J_{soc}$, then for any $i=1,\ldots,N$, $\check{v}_{t}^{i}$ is the solution for minimizing control problem (4.1)-(4.2) when other agents’ maintain their control $\check{v}^{-i}$.
@@
For $i=1,\ldots,N$, let $J_{soc}=Y_{1}^{i}+Y_{2}^{i}$, where
@@
Note that the price process $P$ is only partially influenced by strategy $v^{i}$. Solving (2.2), we obtain
@@
Using the decomposition $Q_{t}=\frac{1}{N}v_{t}^{i}+\frac{1}{N}\sum_{k\neq i}v_{t}^{k}$, we further decompose $Y_{2}^{i}=Y_{3}^{i}+Y_{4}^{i}$, where the components are given by
@@
We thus observe that $J_{soc}=Y_{1}^{i}+Y_{3}^{i}+Y_{4}^{i}$, and both $Y_{1}^{i}$ and $Y_{3}^{i}$ are independent of $v^{i}$. Since $Y_{4}^{i}=\frac{1}{N}\check{J}_{i}(v^{i})$, we conclude that if $(\check{v}_{t}^{1},...,\check{v}_{t}^{N})$ minimizes $J_{soc}$, then for $i=1,\ldots,N$, the strategy $\check{v}_{t}^{i}$ minimizes $\check{J}_{i}(v^{i})$. The proof is complete. ∎
Subsequently, we solve the optimal control problem (4.4). From (4.2), we obtain
@@
Therefore, we consider the system
$\left\{\begin{aligned} &dX_{s}=v_{s}ds+\sigma_{i}dW_{s},\quad X_{t}=x,\\ &du_{s}=(-\alpha u_{s}+\alpha v_{s})ds,\quad u_{t}=u,\quad\forall t\leq s\leq T,\end{aligned}\right$. (4.5)
and we define the admissible strategy set for $v$ as
[…]
@@
$V_{i}(t,x,u)=\inf_{v^{i}\in\mathcal{C}_{i}}\mathbb{E}\left[\int_{t}^{T}\Big(\frac{\eta}{2}(X_{s}^{i}-\kappa)^{2}+\frac{c}{2}(v_{s}^{i})^{2}-\bar{q}(s)u_{s}^{i}+v_{s}^{i}\bar{p}(s)\Big)ds+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
@@
$\frac{\partial V_{i}}{\partial t}+\inf_{v^{i}\in\mathcal{C}_{i}}\left\{\frac{\partial V_{i}}{\partial x}v_{t}^{i}+\frac{\partial V_{i}}{\partial u}(\alpha v_{t}^{i}-\alpha u)+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)\right\}-\bar{q}(t)u+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}V_{i}}{\partial^{2}x}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
with terminal condition $V_{i}(T,x,u)=\frac{\gamma}{2}(x-\zeta)^{2}$. We adopt the ansatz
$V_{i}(t,x,u)=a_{i}(t)x^{2}+b_{i}(t)x+l_{i}(t)u+f_{i}(t),$
which reduces the HJB equation to
$a^{\prime}_{i}(t)x^{2}+b^{\prime}_{i}(t)x+f_{i}^{\prime}(t)+l^{\prime}_{i}(t)u+\inf_{v^{i}\in\mathcal{A}}\Big\{(2a_{i}(t)x+b_{i}(t))v_{t}^{i}+\alpha l_{i}(t)v_{t}^{i}-\alpha ul_{i}(t)$